# Continuous Family of Conformal Field Theories and Exactly Marginal Operators

### v1

@@ IV Examples
where $R$ represents the compactification radius. The marginal operator in this CFT is $\phi={\partial}X\bar{{\partial}}X$. The exactly marginal deformation leads to the shift in the compactification radius, i.e., $R\to R^{\prime}\equiv R\mathrm{e}^{\pi\delta\lambda}$. The interface with a CFT deformed by $\delta\lambda$ can be described by the following gluing matrix [Bachas2002],
$\left(\begin{array}[]{c}{\partial}X_{L}\\ \bar{{\partial}}X_{R}\\ \end{array}\right)=S\left(\begin{array}[]{c}\bar{{\partial}}X_{L}\\ {\partial}X_{R}\\ \end{array}\right),\ \ \ \ \ \ \ S=-\left(\begin{array}[]{cc}\cos 2\theta&-\sin 2\theta\\ \sin 2\theta&\cos 2\theta\\ \end{array}\right),$ (28)
where we define $\tan\theta=\mathrm{e}^{\pi\delta\lambda}$. This gluing condition allows us to express the stress tensor on the right side of the interface in terms of the free boson fields on the left side,
$T_{R}$ $=\frac{1}{2}{\partial}X_{R}{\partial}X_{R}=\frac{1}{2}{\partial}X_{L}{\partial}X_{L}-\pi\delta\lambda{\partial}X_{L}\bar{{\partial}}X_{L},$ (29)
$\overline{T}_{R}$ $=\frac{1}{2}\bar{{\partial}}X_{R}\bar{{\partial}}X_{R}=\frac{1}{2}\bar{{\partial}}X_{L}\bar{{\partial}}X_{L}-\pi\delta\lambda{\partial}X_{L}\bar{{\partial}}X_{L}.$
@@ V Comments
In the introduction, we noted the interesting analogy between the absence of free parameters and the absence of global symmetry in AdS gravitational theories, both of which are consequences of the splittability of symmetry generators. The absence of free parameters has been quantified as Distance Conjectures 1 and 2 [Ooguri:2006in] and formulated in AdS [Perlmutter:2020buo], and some of these conjectures have been proven [Baume:2023msm, Ooguri:2024ofs]. The absence of global symmetry has been quantified as the Weak Gravity Conjecture [Arkani-Hamed:2006emk] and formulated in AdS [Nakayama:2015hga, Harlow:2015lma, Harlow:2018tng, Aharony:2021mpc], and some of these have been proven [Sharon:2023drx, Aharony:2023ike]. Since both are formulated in terms of inequalities, it would be interesting to determine whether they are analogous and can be understood within a single framework.
@@ Acknowledgments
We first compute the three-point function of the normalized displacement operator $\hat{\rm D}\equiv{\rm D}/\sqrt{{\rm N}_{\rm D}}$. For this purpose, we consider the OPE of $T_{L}-T_{R}$,
$\begin{aligned} &(T_{L}(x)-T_{R}(x))(T_{L}(y)-T_{R}(y))(T_{L}(z)-T_{R}(z))\\ &=(\bar{T}_{L}(x)-\bar{T}_{R}(x))(T_{L}(y)-T_{R}(y))(T_{L}(z)-T_{R}(z))\\ &=\bar{T}_{L}(x)T_{L}(y)T_{L}(z)-\bar{T}_{R}(x)T_{R}(y)T_{R}(z)\\ &=\frac{c_{L}\bar{T}_{L}(x)-c_{R}\bar{T}_{R}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left(\bar{T}_{L}(x)T_{L}(z)-\bar{T}_{R}(x)T_{R}(z)\right)\\ &=\frac{c_{L}\bar{T}_{L}(x)-c_{R}\bar{T}_{R}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left(T_{L}(x)T_{L}(z)-T_{R}(x)T_{R}(z)\right)\\ &=\frac{c_{L}\bar{T}_{L}(x)-c_{R}\bar{T}_{R}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left[\frac{c_{L}-c_{R}}{2(x-z)^{4}}+\left(\frac{2}{(x-z)^{2}}+\frac{\partial_{z}}{(x-z)}\right)\left(T_{L}(z)-T_{R}(z)\right)\right].\end{aligned}$ (31)
@@
$\langle(T_{L}(x)-T_{R}(x))(T_{L}(y)-T_{R}(y))(T_{L}(z)-T_{R}(z))\rangle=\frac{c_{L}-c_{R}}{(y-z)^{2}(x-z)^{4}}\left(1+O((y-z))+O((x-z))\right)$ (32)
@@
$c_{\hat{\rm D}\hat{\rm D}\hat{\rm D}}=\frac{c_{L}-c_{R}}{{\rm N}_{\rm D}^{3/2}}\,.$ (33)
Thus, the smoothness assumption implies that $\left|c_{L}-c_{R}\right|^{2/3}/{\rm N}_{\mathrm{D}}$ is finite in the limit ${\rm N}_{\rm D}\to 0$.
@@ V.2 Interface conformal manifolds and $c_{\rho\rho\rho}$
where $\rho$ is a dimension one scalar as in the main text. Here we added an $i$ to make the displacement real for real $\rho$. In this appendix, we prove that the three-point function of $\rho$ vanishes. While the argument starting with equations (21,22) can be repeated with $O_{L}=O_{R}=\rho$, this is to no avail because $\braket{\rho(z,\bar{z})\rho(\bar{z},z)(\partial-\bar{\partial})\rho(w,w)}$ vanishes kinematically. Instead, we will derive the result as a consequence of conformality of the interface, together with the Ward identities of the displacement. In the rest of this appendix, we use the real coordinates $x=\Re z,\,y=\Im z$, and write $\rho(x,y)$ slightly redefining our notation.
@@
$g^{(2)}=g^{(2)}_{R}(\mu)+\frac{5}{16}\,c_{\rho\rho\rho}\,\log a\mu~,\quad\qquad\kappa=\frac{3}{5}~.$ (42)
@@
$\mu\frac{\partial g_{R}^{(2)}}{\partial\mu}=\beta(g_{R}^{(2)})=-\frac{5}{16}\,c_{\rho\rho\rho}~.$ (43)

### v2

@@ IV Examples
where $R$ represents the compactification radius. The marginal operator in this CFT is $\phi={\partial}X\bar{{\partial}}X$. The exactly marginal deformation leads to the shift in the compactification radius, i.e. $R\to R^{\prime}\equiv R\mathrm{e}^{\pi\delta\lambda}$. The interface with a CFT deformed by $\delta\lambda$ can be described by the following gluing matrix [5],
$\left(\begin{array}[]{c}{\partial}X_{1}\\ \bar{{\partial}}X_{2}\\ \end{array}\right)=S\left(\begin{array}[]{c}\bar{{\partial}}X_{1}\\ {\partial}X_{2}\\ \end{array}\right),\ \ \ \ \ \ \ S=-\left(\begin{array}[]{cc}\cos 2\theta&-\sin 2\theta\\ \sin 2\theta&\cos 2\theta\\ \end{array}\right),$ (28)
where we define $\tan\theta=\mathrm{e}^{\pi\delta\lambda}$. This gluing condition allows us to express the stress tensor on the bottom side of the interface in terms of the free boson fields on the upper side,
$T_{2}$ $=\frac{1}{2}{\partial}X_{2}{\partial}X_{2}=\frac{1}{2}{\partial}X_{1}{\partial}X_{1}-\pi\delta\lambda{\partial}X_{1}\bar{{\partial}}X_{1},$ (29)
$\overline{T}_{2}$ $=\frac{1}{2}\bar{{\partial}}X_{2}\bar{{\partial}}X_{2}=\frac{1}{2}\bar{{\partial}}X_{1}\bar{{\partial}}X_{1}-\pi\delta\lambda{\partial}X_{1}\bar{{\partial}}X_{1}.$
@@ V Comments
In the introduction, we noted the interesting analogy between the absence of free parameters and the absence of global symmetry in AdS gravitational theories, both of which are consequences of the splittability of symmetry generators. The absence of free parameters has been quantified as Distance Conjectures 1 and 2 [18] and formulated in AdS [23], and some of these conjectures have been proven [7, 19]. The absence of global symmetry has been quantified as the Weak Gravity Conjecture [4] and formulated in AdS [17, 14, 13, 2], and some of these have been proven [28, 1]. Since both are formulated in terms of inequalities, it would be interesting to determine whether they are analogous and can be understood within a single framework.
@@ Acknowledgments
We first compute the three-point function of the normalized displacement operator $\hat{\rm D}\equiv{\rm D}/\sqrt{{\rm N}_{\rm D}}$. For this purpose, we consider the OPE of $T_{1}-T_{2}$,
$\begin{aligned} &(T_{1}(x)-T_{2}(x))(T_{1}(y)-T_{2}(y))(T_{1}(z)-T_{2}(z))\\ &=(\bar{T}_{1}(x)-\bar{T}_{2}(x))(T_{1}(y)-T_{2}(y))(T_{1}(z)-T_{2}(z))\\ &=\bar{T}_{1}(x)T_{1}(y)T_{1}(z)-\bar{T}_{2}(x)T_{2}(y)T_{2}(z)\\ &=\frac{c_{1}\bar{T}_{1}(x)-c_{2}\bar{T}_{2}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left(\bar{T}_{1}(x)T_{1}(z)-\bar{T}_{2}(x)T_{2}(z)\right)\\ &=\frac{c_{1}\bar{T}_{1}(x)-c_{2}\bar{T}_{2}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left(T_{1}(x)T_{1}(z)-T_{2}(x)T_{2}(z)\right)\\ &=\frac{c_{1}\bar{T}_{1}(x)-c_{2}\bar{T}_{2}(x)}{2(y-z)^{4}}+\left(\frac{2}{(y-z)^{2}}+\frac{\partial_{z}}{(y-z)}\right)\left[\frac{c_{1}-c_{2}}{2(x-z)^{4}}+\left(\frac{2}{(x-z)^{2}}+\frac{\partial_{z}}{(x-z)}\right)\left(T_{1}(z)-T_{2}(z)\right)\right].\end{aligned}$ (31)
@@
$\langle(T_{1}(x)-T_{2}(x))(T_{1}(y)-T_{2}(y))(T_{1}(z)-T_{2}(z))\rangle=\frac{c_{1}-c_{2}}{(y-z)^{2}(x-z)^{4}}\left(1+O((y-z))+O((x-z))\right)$ (32)
@@
$c_{\hat{\rm D}\hat{\rm D}\hat{\rm D}}=\frac{c_{1}-c_{2}}{{\rm N}_{\rm D}^{3/2}}\,.$ (33)
Thus, the smoothness assumption implies that $\left|c_{1}-c_{2}\right|^{2/3}/{\rm N}_{\mathrm{D}}$ is finite in the limit ${\rm N}_{\rm D}\to 0$.
@@ V.2 Interface conformal manifolds and $c_{\rho\rho\rho}$
where $\rho$ is a dimension one scalar as in the main text. Here we added an $i$ to make the displacement real for real $\rho$. In this appendix, we prove that the three-point function of $\rho$ vanishes. While the argument starting with equations (21,22) can be repeated with $O_{1}=O_{2}=\rho$, this is to no avail because $\braket{\rho(z,\bar{z})\rho(\bar{z},z)(\partial-\bar{\partial})\rho(w,w)}$ vanishes kinematically. Instead, we will derive the result as a consequence of conformality of the interface, together with the Ward identities of the displacement. In the rest of this appendix, we use the real coordinates $x=\Re z,\,y=\Im z$, and write $\rho(x,y)$ slightly redefining our notation.
@@
$g^{(2)}=g^{(2)}_{2}(\mu)+\frac{5}{16}\,c_{\rho\rho\rho}\,\log a\mu~,\quad\qquad\kappa=\frac{3}{5}~.$ (42)
@@
$\mu\frac{\partial g_{2}^{(2)}}{\partial\mu}=\beta(g_{2}^{(2)})=-\frac{5}{16}\,c_{\rho\rho\rho}~.$ (43)